OpenStudy (anonymous):

simplify (1+csc x)/cos x+cot x

OpenStudy (anonymous):

(1 + csc x) / (cos x + cot x) = sin x (1 + csc x) / sin x (cos x + cot x) = sin x + 1 / (sin x cos x + cos x) = sin x + 1 / cos x (sin x + 1) = 1 / cos x or just sec x so you were correct :)

OpenStudy (abb0t):

How do you know how to simplify this @Keandrap15 but couldn't factor out??

OpenStudy (anonymous):

My sister did it when i was gone from my computer.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome